OpenStudy (girlstudy):

An object starts from rest with a constant acceleration of 8.00 m/s along a straight line. Find (a) the speed at the end of 5.00 s, (b) the average for the 5¬s interval, and (c) the distance traveled in the 5.00 s.

OpenStudy (jiteshmeghwal9):

acceleration(a)=8 m/s^2 initial velocity =0 a).Here we can apply equation of motion as object is moving with constant acceleration v=u+at v=final speed a= acceleration t=time u=initial speed

OpenStudy (jiteshmeghwal9):

c). as we know object is moving with constant acceleration so \(s= ut+\frac{1}{2}at^2\) s=distance travelled u=initial speed a=acceleration t=time

OpenStudy (jiteshmeghwal9):

I just don't gt wht r u asking in part b. I mean that average of which quantity do u want to find ?